OpenStudy (aakashsudhakar):

Can someone help with this integration? We just learned integrations by manipulation of quadratic expressions. ∫(1/((x^2 + 4x + 5)^2))dx

OpenStudy (aakashsudhakar):

This may help with visualizing it. \[∫\frac{ dx }{ (x^2 + 4x + 5)^2 }\]

OpenStudy (amistre64):

by manipulation of quadratics eh .. never heard of that method. Can you explain it?

OpenStudy (amistre64):

the wolf gives a nice result, so the process would seem to be relatively easy to follow; im just not aware of the method

OpenStudy (aakashsudhakar):

Essentially, from how I understood it, you get the denominator into a square form by adding constants inside and outside of the expression: ((x^2 + 4x + 4) - 4 + 5)^2 ((x + 2)^2 + 1)^2 I thought you could use u-substitution here, but it doesn't get me anywhere (at least I don't think it does)... u = x + 2 du = dx --> ((u^2) + 1)^2 So now you'd have... \[∫\frac{ du }{ (u^2 + 1)^2 }\] It's at this step that I'm confused as to what to do next. I have a feeling that I have to do some wacky inverse trig substitutions, but I can't identify what exactly I have to do...

OpenStudy (amistre64):

from there it would be partial decomp ... \[\frac{ 1 }{ (u^2 + 1)^2 }=\frac{Au+B}{u^2+1}+\frac{Cu+D}{(u^2+1)^2}\]

OpenStudy (amistre64):

when u=0 1 = B + D when u = 1 1 = (A + B )/2 + (C+D)/4 4 = 2A + 2B + C + D if we develop 4 equations in 4 unknowns, we can solve for ABCD

OpenStudy (amistre64):

i did bad when u=1 .. i ignored the denominator on the left :( hard part is keeping track is all

OpenStudy (amistre64):

when u = 1 1/4 = (A + B )/2 + (C+D)/4 1 = 2A + 2B + C + D thats better

OpenStudy (amistre64):

when u = -1 1/4 = (-A + B )/2 + (-C+D)/4 1 = -2A + 2B - C + D for simplicity, let u = 2 1/25 = (2A + B )/5 + (2C+D)/25 1 = 10A + 5B + 2C + D 0A + 1B + 0C+ 1D = 1 2A + 2B + 1C + 1D = 1 -2A + 2B - 1C + 1D = 1 10A + 5B + 2C + 1D = 1 well that didnt turn out as well as i was hoping it would :/ ABC = 0, D = 1

OpenStudy (aakashsudhakar):

I was actually gonna say... I think I managed to work out a way to fit in a u-substitution. Could you say that, if going back to... \[∫\frac{ du }{ (u^2 + 1)^2 }\] u = tan ø du = sec^2(ø) dø ...and then work from there?

OpenStudy (amistre64):

its worth a shot

OpenStudy (amistre64):

id love to stay and work over it but my times up, going home for the night ....

OpenStudy (aakashsudhakar):

NP, thanks anyway! I think I can work my way through it eventually...

OpenStudy (anonymous):

Yes, you're right about the trig sub: \[\int\frac{ du }{ (u^2 + 1)^2 }\] \(u=\tan t\), so \(du=\sec^2t ~dt\): \[\int\frac{ \sec^2t }{ (\tan^2t + 1)^2 }~dt\\ \int\frac{ \sec^2t }{ (\sec^2t)^2 }~dt\\ \int\frac{dt}{ \sec^2t}\\ \int \cos^2t~dt\]

OpenStudy (anonymous):

@amistre64, not sure, but i think the partial fraction decomp will result in the same function, i.e. A=B=C=0 and D=1.

OpenStudy (aakashsudhakar):

From ∫cos^2(t) dt... \[\frac{ 1 }{ 2 }∫(1 + \cos 2\theta) d \theta \] = \[\frac{ 1 }{ 2 }∫d \theta + \frac{ 1 }{ 2 }∫ \cos 2\theta d \theta\] = \[\frac{ \theta }{ 2 } + \frac{ \sin 2\theta }{ 4 } + C\] And then just convert theta Ø back into terms of X, and that's it?

OpenStudy (anonymous):

Yep, that's correct

OpenStudy (amistre64):

yeah, when I rrefed the matrix i was horrified!! lol